A farmer has 240 meters of fencing and wants to enclose a rectangular area of 3200 square meters. What dimensions should he use?
Answer by Alan3354(69443)

A farmer has 240 meters of fencing and wants to enclose a rectangular area of 3200 square meters. What dimensions should he use?
----------------------
L*W = 3200
2L + 2W = 240 --> W = 120-L
--------------
L*(120-L) = 3200
120L - L^2 = 3200
L^2 - 120L + 3200 = 0
(L-40)*(L-80) = 0
L = 40, W = 80
L = 80, W = 40